luhivqqcherry:

helpp with mathh againn šŸ¤ ! . A typical cell phone plan has a fixed base fee that includes a certain amount of data and an overage charge for data use beyond the plan. A cell phone plan charges a base fee of $62 and an overage charge of $30 per gigabyte of data that exceed 2 gigabytes. If C represents the cost and g represents the total number of gigabytes of data, which equation could represent this plan when more than 2 gigabytes are used? a) C=30+62(2āˆ’g) b) C=30+62(gāˆ’2) c) C=62+30(gāˆ’2) d) C=62+30(2āˆ’g)

Phantomdex:

I think your answer would be c = 62 + 30 ( g - 2) it seems (to me) that it would be the one that aligns most with the equation.

Yourstrulyfr:

c) C = 62 + 30(g - 2) Explanation: The base fee of $62 is included regardless of the amount of data used, so it is added to the equation. For data beyond 2 gigabytes, there is an overage charge of $30 per gigabyte. , (g - 2) represents the excess gigabytes beyond the 2-gigabyte limit.

luhivqqcherry:

@phantomdex wrote:
I think your answer would be c = 62 + 30 ( g - 2) it seems (to me) that it would be the one that aligns most with the equation.
@yourstrulyfr wrote:
c) C = 62 + 30(g - 2) Explanation: The base fee of $62 is included regardless of the amount of data used, so it is added to the equation. For data beyond 2 gigabytes, there is an overage charge of $30 per gigabyte. , (g - 2) represents the excess gigabytes beyond the 2-gigabyte limit.
ahh tysmm bothh of youu !! .

Phantomdex:

Also nice pfp cherry :>

luhivqqcherry:

@phantomdex wrote:
Also nice pfp cherry :>
aww tysmm !! .

Yourstrulyfr:

Yww!!

Phantomdex:

@luhivqqcherry wrote:
@phantomdex wrote:
I think your answer would be c = 62 + 30 ( g - 2) it seems (to me) that it would be the one that aligns most with the equation.
@yourstrulyfr wrote:
c) C = 62 + 30(g - 2) Explanation: The base fee of $62 is included regardless of the amount of data used, so it is added to the equation. For data beyond 2 gigabytes, there is an overage charge of $30 per gigabyte. , (g - 2) represents the excess gigabytes beyond the 2-gigabyte limit.
ahh tysmm bothh of youu !! .
Ofc, I may not be the best at math but I'll help when I can!! ^^